OpenStudy (bigbosssaint21):

the length of two sides of a triangle are, side 1 , 8x^2-5x-2, side 2 ,7x-x^2+3 the perimeter of the triangle is 4x^3-3x^2+2x-6 what is the length of the third side of the triangle?

ganeshie8 (ganeshie8):

say the length of third side is \(\large a\)

ganeshie8 (ganeshie8):

since perimeter is the sum of all the side lengths : \(\large a + (8x^2-5x-2 ) + ( 7x-x^2+3 ) = 4x^3-3x^2+2x-6 \)

ganeshie8 (ganeshie8):

isolate \(\large a\)

OpenStudy (bigbosssaint21):

how?

ganeshie8 (ganeshie8):

group like terms : put all x^2 terms together, x terms together constants together

ganeshie8 (ganeshie8):

\(\large a + (8x^2 - x^2) + (-5+7x) + (-2+3) = 4x^3-3x^2+2x-6 \)

ganeshie8 (ganeshie8):

simplifying a bit gives you : \(\large a + (7x^2) + (2x) + (1) = 4x^3-3x^2+2x-6 \)

ganeshie8 (ganeshie8):

are you with me so far ? :)